Gould Corporation uses the following activity rates from its activity-based costing to assign overhead costs to products:

Activity Cost Pool Activity Rate
Setting up batches $ 59.91 per batch
Processing customer orders $ 73.17 per customer order
Assembling products $ 4.60 per assembly hour

Data concerning two products appear below:

Product K91B Product F65O
Number of batches 96 67
Number of customer orders 46 60
Number of assembly hours 500 907

How much overhead cost would be assigned to Product K91B using the activity-based costing system?

Multiple Choice

  • $5,751.36

  • $137.68

  • $102,754.62

  • $11,417.18
Answer #1

Answer

$11,417.18

Calculations:

Cost
Setting up batches $     5,751.36
Processing customer orders $     3,365.82
Assembling products $     2,300.00
Total $ 11,417.18

Note:

Cost
Setting up batches =59.91*96
Processing customer orders =73.17*46
Assembling products =4.6*500

in case of any doubt, please comment.